OpenStudy (shawn):

if the domain of sin(x) is restricted to [-2pi,2pi], is it still a periodic function

OpenStudy (loser66):

nope,

OpenStudy (shawn):

why not? there are two cycles right?

OpenStudy (loser66):

oh, yeah, you are right. My stupid comment. :)

OpenStudy (shawn):

but I thought it's not periodic just like you did though :/

OpenStudy (shawn):

despite there are two cycles

OpenStudy (loser66):

2 is enough!!

OpenStudy (shawn):

well, i'm still confused about the definition. def: A function is said to be periodic if there exists a non-zero real number P such that f(x+P) = f(x) for all x in the domain.

OpenStudy (shawn):

so if the period is 2pi (which it is), the value of sin(2pi + 2pi) is undefined , so we can't say sin(2pi + 2pi) = sin(2pi)

OpenStudy (shawn):

somehow the definition implies there are infinitely many cycles

OpenStudy (loser66):

Hey!! why not? sin 4pi = sin 2pi =0

OpenStudy (loser66):

who says sin (4pi) is undefined?

OpenStudy (shawn):

well, because 4pi isn't in the domain [-2pi,2pi]

OpenStudy (loser66):

out of domain?

OpenStudy (loser66):

it doesn't mean sin(4pi) is undefined, to me!!

jimthompson5910 (jim_thompson5910):

If you didn't have the domain restriction, then yes it would be periodic. The domain restriction makes the function not be periodic.

OpenStudy (shawn):

@Loser66, well sin(4pi) is undefined. If you draw a vertical line at x = 4pi, it doesn't intersect the function. @jim_thompson5910 I though so too.

OpenStudy (loser66):

@jim_thompson5910 I don't think so. It is still http://mathworld.wolfram.com/PeriodicFunction.html

OpenStudy (loser66):

it is doubly periodic function

OpenStudy (shawn):

@jim_thompson5910 so somehow being period implies there are infinitely many cycles, which I need someone to confirm that. if it is not periodic, then the negation of the definition is: For all non-zero value P, there exists a element x in the domain such that f(x + P) ≠ f(x)

OpenStudy (shawn):

"being periodic"

jimthompson5910 (jim_thompson5910):

Then again I can't find any issues if I plug in values from the domain. So it appears sin(x) is periodic

jimthompson5910 (jim_thompson5910):

x = 4pi isn't in the domain of [-2pi,2pi] but we don't plug in x = 4pi x = 2pi is the value being plugged in which is in the domain

OpenStudy (shawn):

remember that the domain is [-2pi,2pi] instead of (-inf,inf)

OpenStudy (shawn):

right. We plug in x = 2pi but we still need to value of sin(4pi) right?

OpenStudy (shawn):

well, let's just take a look at what it means to be not periodic. Any non-zero value P besides P = 2pi will satisfy the statement. for example if P = pi, then choose x = pi/2 sin(pi/2 + pi) ≠ sin(pi/2). But if P = 2pi, that what value of x will work? clearly, we are back the same issue of being out of the domain

OpenStudy (shawn):

doex x = 2pi work? can I say sin(2pi + 2pi) ≠ sin(pi) ? If this is true, then the function is not periodic

OpenStudy (loser66):

Even the domain is [-2pi,2pi], It doesn't mean the range of the function is undefined. And the range of sin (x) is [-1,1] for all x. And that range make the function periodic, not the domain, to me

OpenStudy (shawn):

sorry, sin(2pi + 2pi) ≠ sin(2pi)

OpenStudy (shawn):